OpenStudy (anonymous):

Write the number in scientific notation. In a certain city, the subway system carried a total of 1,890,000,000 passengers

OpenStudy (anonymous):

1.89x10^9, just count the number of decimal places

OpenStudy (anonymous):

10^1=10 10^2=100 10^3=1000 10^4=10000 10^5=100000 ... 1,000,000,000=10^9 1890,000,000=1.89*10^9